Mary’s age is 8years and her father is 43 years, In how many year’s time will the father be 3 times as old as his daughter. Please, show the steps. Thanks.

Answer by ankor@dixie-net.com(22740)

let y = no. of years for this to be true
:
" Mary’s age is 8years and her father is 43 years,
In how many year’s time will the father be 3 times as old as his daughter."
y + 43 = 3(y + 8)
distribute the 3
y + 43 = 3y + 24
subtract y from both sides
43 = 3y - y + 24
subtract 24 from both sides
43 - 24 = 2y
19 = 2y
divide both sides by 2
y = 19/2
y = 9.5 yrs
:
Usually in these problems you get an integer solution but let's see if this is true.
Their ages in 9.5 yrs
43 + 9.5 = 52.5 yrs is Dad's age
8 + 9.5 = 17.5 is Mary's age
Check: 3 * 17.5 = 52.5